OpenStudy (anonymous):

find dr/d(theta): theta^1/2+r^1/2=1

OpenStudy (harsimran_hs4):

differentiate wrt to theta the whole equation can you proceed ?

OpenStudy (anonymous):

what is wrt

OpenStudy (harsimran_hs4):

wrt = with respect to i think i should have written it as w.r.t

OpenStudy (anonymous):

1/2theta^-1/2+1/2r^-1/2

OpenStudy (harsimran_hs4):

alright there are little mistakes given a function f(r) = r tell me what is the derivative of f(r) w.r.t theta?

OpenStudy (anonymous):

does theta represent x and r represent y

OpenStudy (harsimran_hs4):

just think it this way that there are two variables r and theta if you really want x and y then say f(y) = y tell me derivative of f(y) w.r.t x?

OpenStudy (anonymous):

1/2y^-1/2 dy/dx

OpenStudy (harsimran_hs4):

why 1/2 y^-1/2 ??? what is derivative of y wrt y ?

OpenStudy (anonymous):

when substracting 1/2-1 =-1/2

OpenStudy (harsimran_hs4):

derivative of y wrt y is 1 i think isn`t it?

OpenStudy (anonymous):

i believe so

OpenStudy (anonymous):

yes

OpenStudy (harsimran_hs4):

then from where did 1/2 come in ?

OpenStudy (anonymous):

when doing the derivative you have to move the exponent in front of the variable then subtract the exponent from one so... you have y^1/2 the derivative would be 1/2y^-1/2

OpenStudy (anonymous):

because 1/2-1/1=-1/2

OpenStudy (harsimran_hs4):

that is fine but i never asked for derivative of y^1/2 , i asked you for derivative of y

OpenStudy (anonymous):

1

OpenStudy (harsimran_hs4):

alright now switch back to the above eq (in the question) and differentiate it wrt to theta and tell me what you get?

OpenStudy (anonymous):

y represents theta for this equation

OpenStudy (harsimran_hs4):

I don`t know what the confusion is all about earlier you had x and y and the two variable now it`s r and theta the steps to solve still remain the same

OpenStudy (callisto):

Find \(\frac{dr}{d\theta}: \) for \[ \theta^\frac{1}{2}+r^\frac{1}{2}=1\] If you are more comfortable with x and y, then consider our usual practice, we just let r=y and x=θ here. So, we get \[ x^\frac{1}{2}+y^\frac{1}{2}=1\] Now, differentiate both sides with respect to x. What do you get?

OpenStudy (anonymous):

1/2theta^-1/2+1/2r^-1/2dr/dtheta (1/1/2theta^1/2)+(1/1/2r^1/2dr/dtheta)=0

OpenStudy (callisto):

The first step looks fine, except there should be an "=0" after dr/d theta \[\frac{1}{2}\theta^{-\frac{1}{2}}+\frac{1}{2}r^{-\frac{1}{2}}\frac{dr}{d\theta}=0\] For the second step, we need to use the property that \(\large a^{-x}=\frac{1}{a^x}\). It should be like this: \[\frac{1}{2\theta^{\frac{1}{2}}}+\frac{1}{2r^{\frac{1}{2}}}\frac{dr}{d\theta}=0\]

OpenStudy (anonymous):

why isn't it 1/1/2(theta)^1/2 if 1/2 is the number in front of the variable

OpenStudy (callisto):

\[\frac{1}{2}\theta^{-\frac{1}{2}}=\frac{1}{2}\times \frac{1}{\theta^{\frac{1}{2}}}=\frac{1}{2\theta^{\frac{1}{2}}}\]

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

\[1/\sqrt{2\theta}+1/\sqrt{2r}\]

OpenStudy (anonymous):

well technically \[1/\sqrt{2\theta}+1/\sqrt{2r}dy/dx\]

OpenStudy (anonymous):

my guess is to add the fractions

OpenStudy (callisto):

No?! No square root for 2s.

OpenStudy (anonymous):

so it would just be \[1/2 \theta ^{1/2} +1/2r ^{1/2}dr/d \theta \]

OpenStudy (anonymous):

then add the fractions

OpenStudy (callisto):

Hmmm \[\frac{1}{2\theta^{\frac{1}{2}}}+\frac{1}{2r^{\frac{1}{2}}}\frac{dr}{d\theta}=0\]And then, you should isolate dr/d theta

OpenStudy (anonymous):

\[1/2r ^{1/2} dr/d \theta =-1/2\theta ^{1/2}\]

OpenStudy (anonymous):

then multiply 2r^1/2/1 on both sides i think

OpenStudy (callisto):

(2r^1/2)/1 - Yes :)

OpenStudy (anonymous):

\[-1/2 \theta ^{1/2} (2\theta ^{1/2})=-2\theta ^{1/2}/2\theta ^{1/2}=-1\]

OpenStudy (anonymous):

ooops wrong

OpenStudy (anonymous):

\[-2r ^{1/2}/2\theta ^{1/2}=-r ^{1/2}/\theta ^{1/2}=-\sqrt{r}/\sqrt{\theta}\]

OpenStudy (callisto):

\[\frac{1}{2\theta^{\frac{1}{2}}}+\frac{1}{2r^{\frac{1}{2}}}\frac{dr}{d\theta}=0\]\[\frac{1}{2r^{\frac{1}{2}}}\frac{dr}{d\theta}=-\frac{1}{2\theta^{\frac{1}{2}}}\]\[\frac{2r^{\frac{1}{2}}}{1}\times \frac{1}{2r^{\frac{1}{2}}}\frac{dr}{d\theta}=-\frac{2r^{\frac{1}{2}}}{1}\times \frac{1}{2\theta^{\frac{1}{2}}}\]\[\frac{dr}{d\theta}=\frac{-2r^{\frac{1}{2}}}{2\theta^{\frac{1}{2}}}=...\] Yes!!~

OpenStudy (anonymous):

Thank you again =^^=

OpenStudy (callisto):

Welcome :)